OpenStudy (anonymous):

What is the common ratio of the geometric sequence 16, 24, 36, 54, ...? 2/3 3/2 1 8

OpenStudy (the_fizicx99):

Divide, 54/36 = 1.5 36/24 = 1.5 24/16 = 1.5 The average ratio is 1.5 or 3/2